What are the four basic arithmetic operators?

Give meaningful example of using each.

Here is an example with the remainder operator:

Remainder Operator: %, returns the remainder after division, 7%3 = 1.

piecesOfCandyAfterEqualDivision = piecesOfCandy % numOfPeople
Answer #1

An arithmetic operator takes two operands and performs the particular operation on them.

The four basic arithmetic operators are :

Addition (+)

Subtraction (-)

Multiplication (*)

Division (/)

Let’s see all the examples :

+ (Addition)

If the distance from A to B is 5 kms and distance from B to C is 10 kms ,

Then the distance from A to C = Dist. From A to B + Dist. From B to C

- In a programming language say C ,

We write is as , x= y+z;

- (Subtraction)

If bob has 5 oranges and he eats 2 oranges from them.

Number of oranges left = 5- 2 = 3

- In a programming language say C ,

We write is as , x= y-z;

* (Multiplication)

If 2 cats have 5 kittens each , then total number of kittens = 5*2 =10

- In a programming language say C ,

We write is as , x= y*z ;

/ (Division)

If we have 10 apples and we have to give it 2 people , then

Number of apples each person will have = 10/2 = 5 apples each

- In a programming language say C ,

We write is as , x= y/z;